OpenStudy (anonymous):

2sin2theta=1/2, find the value of theta

OpenStudy (sidsiddhartha):

use first sin2(theta)=2sin(theta)cos(theta) then use euler's formula

OpenStudy (sidsiddhartha):

\[cosx=\frac{ e ^{ix}+e ^{-ix} }{ 2 }\] \[sinx=\frac{ e ^{ix}-e ^{-ix} }{ 2i}\] use them

OpenStudy (john_es):

\[2\sin (2\theta)=1/2\Rightarrow \sin (2\theta)=1/4\Rightarrow\\ \Rightarrow 2\theta=\arcsin(1/4)\Rightarrow \theta=\frac{1}{2}\arcsin(1/4)\]